# A Single Graph Convolution is All You Need: Efficient Grayscale Image Classification

### v3

@@ 1 Introduction
With the recent technological advances, modern FPGAs contain many hardware resources [2], including DSPs, LUTs, BRAMs, and URAMs. The programmable nature of FPGAs allows users to develop a customized data path and on-chip memory organization, leading to high-performance implementations. Consequently, FPGAs have emerged as an appealing option for executing time-sensitive machine learning tasks with reduced latency and power. For instance, FPGAs have been used for accelerating machine learning [3] and graph analytic tasks [4]. Given that our model is a lightweight image classifier, we find it suitable to be deployed on an FPGA. We can perform inference on the FPGA without returning the intermediate results to external memory. This ensures low-latency inference by capitalizing the fine-grained data parallelism inherent in FPGAs. In contrast, CPUs and GPUs exploit coarse-grained thread-level parallelism and have complex cache hierarchies, which are unsuitable for low-latency inference. Thus, this paper makes the following contributions:
@@
  • We implement our model on a state-of-the-art FPGA board, Xilinx Alveo U200. Compared with state-of-the-art GPU implementation, our FPGA implementation achieves 2.78×2.78$\times$2.78 × speedup in latency and 2.1×2.1$\times$2.1 × speedup in throughput.
@@ 3.1 MLP Approaches
Our model combines various components of simple models and is inherently different from current works in low-latency image classification. Some recent architectures involve simple MLP-based models. Touvron et al. introduced ResMLP [5], an image classifier based solely on MLPs. ResMLP is trained in a self-supervised manner with patches interacting together. Touvron et al. highlight their model’s high throughput properties and accuracy. ResMLP uses patches from the image and alternates linear layers where patches interact and a two-layer feed-forward network where channels interact independently per patch. Additionally, MLP-Mixer [6] uses a similar patching method, which also attains competitive accuracy on RGB image datasets compared to other CNNs and transformer models. Our proposed method uses the results from a single-layer MLP to feed into a graph neural network, during which we skip the information from the three-channel RGB setting and only consider the single-channel grayscale problem. This is inherently different than the methods mentioned earlier, as they use patching approaches while we focus on the vectorization of pixels.
@@ 3.2 Graph Image Construction Methods
Han et al. [10] form a graph from the image by splitting the image into patches, much like a transformer. A deep graph neural network learns on the patches similarly to a transformer but in a graph structure. Our structure does not form a graph where each patch is a node in a graph; instead, we create a graph from the resultant of a vectorized image passed through a single-layer MLP.
@@ 4 Overview and Architecture
Our model vectorizes a batch of images, allowing us to use a fully connected layer pixel-wise for low computation time rather than relying on convolutional neural networks. The proposed method uses the resultant from MLPs and constructs a graph batch-wise, where each node corresponds to the flattened image outputted from an MLP. We then apply a batch-wise attention term, inputted into an MLP for classification. Image Vectorization. For each image in a batch, we view the image as a vector. For a tensor $\mathbf{X}\in\mathbb{R}^{B\times H\times W}$ where $B$ is the batch size, $H$ and $W$ are the height and width of an image, we flatten the tensor to $\mathbf{X}_{1}\in\mathbb{R}^{B\times(H\cdot W)}$. Viewing an image as a vector allows our model to skip the traditional convolutional neural network, which views the image as a grid and cuts computation time.
@@
Simple MLP Layer. We input $\mathbf{X}_{1}$ into a linear MLP with output dimensionality $D_{out}$. Formally, $\text{\scriptsize{MLP}}(\mathbf{X}_{1})\rightarrow\mathbf{X}_{2}\in\mathbb{R}^{B\times D_{out}}$. Formally, MLP = $\sigma(\sum_{j}w_{j}x_{j}+b)$ where $\sigma$ is a general activation term, each $w$ is a learned weight vector, each $x$ is a data vector, and $b$ is a bias term. We then apply a standard layer norm and dropout functions. We use a standard activation function to $\mathbf{X}_{2}$, followed by the standard dropout and MAXPOOL1D functions.
@@
Graph Construction. Our graph construction consists of using each resultant vector from MLP as a node in a graph, such that the number of nodes in the graph corresponds with the number of images in a batch. This means that for each batch, a vectorized image is each node in the graph with feature size $\mathbb{R}^{D_{out}/2}$, and each image is connected to every other image in a batch. Formally, we calculate the adjacency matrix $\mathbf{A}$ as $\mathbf{A}_{ij}=1$, which connects all nodes. We then perform message passing amongst vectors, meaning the message that is aggregated from $i$’s graph neighborhood, given its neighbors $j$.
@@
Graph Convolution. Our single GCN layer learns from similar features of images within its minibatch. Generally, a graph convolutional layer updates the representations of nodes by aggregating each node’s neighbor’s representation. It takes an input $\bm{h_{i}}$ and maps an output $\bm{h_{i}}\mapsto\bm{h^{\prime}_{i}}$ given its neighbors $\mathcal{N}_{i}=\{j\in\mathcal{V}\mid(j,i)\in\mathcal{E}\}$ where $\bm{h^{\prime}}_{i}$ is defined as
$\bm{h^{\prime}}_{i}=f_{\theta}\left(\bm{h}_{i},\text{\scriptsize{AGGREGATE}}\left(\{\bm{h_{j}}\mid j\in\mathcal{N}_{i}\}\right)\right).$
In our case, the input for each node $\bm{h_{i}}$ is the output from the MLP layer of a vectorized image. We claim that this component increases and stabilizes accuracy in the case of our simple MLP-based model. Section 5.2.1 presents a study of accuracy on MNIST with the graph convolutional layer varying. We find that the Residual Gated Graph Convolutional Network [12] outperforms in all categories, and thus, we use this as a backbone for our model.
From the result of inputting $\mathbf{X}_{2}$ into the previous dropout and MAXPOOL1D functions, we fix the resultant matrix $\mathbf{X}_{3}$. From these functions, $\mathbf{X}_{3}$ inputs into a single-layer GCN with dimension $\mathbb{R}^{B\times D_{out}}$ and maps to the same output dimension $\mathbb{R}^{B\times D_{out}}$ where the output is denoted $\mathbf{X}_{4}$.
Thus, we can denote a general GCN term output $\mathbf{X}_{4}$. We then consider the output term layer as $\mathbf{X}_{4}$ as $\mathbf{X}_{4}=\sigma\left(\mathbf{A}\mathbf{X}_{3}\mathbf{W}\right)$ for the general GCN case where $\mathbf{W}$ is a learned layer.
From the resultant matrix $\mathbf{X}_{4}$, we apply a standard batch normalization term, activation function, and MAXPOOL1D functions.
@@
Batch-wise Attention, Residual Connections, & Output. Then, we propose the addition of a residual connection with batch-wise attention, defined as
@@
where $\sigma$ is the standard sigmoid function. The batch-wise attention term allows the model to further capture similar features from each image to another batch-wise. Relating similar properties from images to each other boosts accuracy in our case; refer to our ablation study in Table 7.
The residual term is defined $\mathbf{X}_{6}=\mathbf{X}_{5}+\mathbf{X}_{4}$. The residual term allows our model to capture similar features while minimizing the risk of gradient explosion and retains information from the previous $\mathbf{X}_{4}$ step. By multiplying a softmax-like term with the output of a previous graph convolution, our intuition is that we weigh the correspondence of each pixel compared to other similar pixels within similar images batch-wise. This method differs from the standard query-key-value attention mechanism. It is more suitable to our case, as we allow batch-wise feature assimilation, allowing the model to capture similar features between images in a batch. In contrast, attention mechanisms focus on individual features, which may not be as effective in capturing batch-wise similarities. We then feed the residual term into an MLP for classification results. [^1] We make our code publicly available at https://github.com/GECCOProject/GECCO
@@ 5.1 Datasets
  • MSTAR is a SAR ATR dataset with a training size of $2747$ and testing size of $2425$ SAR images of $10$ different vehicle categories. We resize this dataset to $(128,128)$ pixels.
@@ 5.2.1 Backbone
GCN [16] $97.09\%$ $54.58\pm 8.47$ $4.68\pm 0.72$
TAGConv [17] $97.39\%$ $58.39\pm 9.01$ $4.37\pm 0.67$
SAGEConv [18] $97.77\%$ $58.84\pm 8.98$ $4.32\pm 0.66$
ChebConv [19] $97.50\%$ $51.98\pm 8.65$ $4.91\pm 0.82$
ResConv [12] $\bf{98.04\%}$ $\bf{62.44\pm 9.87}$ $\bf{4.10\pm 0.65}$
@@ 5.2.2 Experimental Performance
In Tables 2, 3, 4, and 5, we present a summary of our findings. We report the best-performing accuracy, average throughputs, and latencies with their standard deviations. Our model outperforms every other model in terms of throughput and latency across all datasets, leads accuracy on the MSTAR dataset, and performs competitively in terms of accuracy on all datasets.
@@ 5.2.3 Model Complexity Metrics
Model complexity metrics for this paper include the number of multiply-accumulate operations, the number of model parameters, model size, and the number of layers. In other words, suppose accumulator $a$ counts an operation of arbitrary $b,c\in\mathbb{R}$. We count the number of multiply-accumulate operations as $a\leftarrow a+(b\times c)$. Additionally, the layer count metric is an essential factor of latency. Increasing the number of layers will also improve the latency of a model’s inference. The goal of an effective machine learning model is to maximize throughput while minimizing the number of MACs and the number of layers, in our case.
@@
GECCO $\bf{1.22\times 10^{6}}$ $\bf{1.90\times 10^{4}}$ $\bf{0.075}$ $\bf{10}$
@@ 5.3 Discussion
Across multiple datasets, GECCO achieves leading or competitive accuracy compared to other state-of-the-art image classifiers. With respect to model complexity, it is clear that GECCO outperforms other machine learning models, highlighting our model’s low latency and lightweight properties.
It is difficult for our model to generalize to the RGB setting. We attribute this challenge to the vectorization process that our model uses. Learning on three channels poses a complexity challenge, as GECCO is very shallow and simple, thus making it challenging to learn on three separate channels.
Our proposed method does not make use of positional embeddings or class tokens. GECCO can learn essential features by the weighted residual term. Additionally, we tested the addition of positional embeddings and class tokens and found no improvement in accuracy across various datasets. We note that the $\mathbf{X}_{5}$ residual term adds positional awareness to the model.
@@ 5.4 FPGA Implementation
We develop an accelerator for the proposed model on a state-of-the-art FPGA, Xilinx Alveo U200 board, to highlight the model’s efficiency and compatibility with hardware further. It has 3 Super Logic Regions (SLRs), 4 DDR memory banks, 1182k Look-up tables, 6840 DSPs, 75.9 Mb of BRAM, and 270 Mb of URAM. The FPGA kernels are developed using the Xilinx High-level Synthesis (HLS) tool to expedite the design process.
We perform the following optimizations for our FPGA design: (1) Portable design: We design a parameterized hardware template using HLS. It is portable to different FPGA platforms, including embedded and data-center FPGAs. (2) Resource sharing: The model is executed layer-by-layer. Each layer in the model is decomposed into basic kernel functions. The basic kernel functions, including matrix multiplication, elementwise activation, column-wise and row-wise summations, max pooling, and various other elementwise operations, are implemented separately and subsequently invoked within their corresponding layers. Due to the reuse of these fundamental kernel functions across multiple layers, FPGA resources are shared among the different layers, maximizing runtime hardware resource utilization. (3) One-load strategy: We employ a one-time data load strategy that enables us to load the required data from DDR only once. All other data required for the computations are stored in on-chip memory, reducing inference latency. Figure 2 illustrates the overall hardware architecture of our design.
@@
We compare our FPGA implementation with the baseline GPU implementation. The GPU baseline is executed on an NVIDIA RTX A5000 GPU, which operates at 1170 MHz and has a memory bandwidth of 768 GB/s. On the other hand, the FPGA operates at 300 MHz and has an external memory bandwidth of 77 GB/s. The GPU baseline is comparable with the FPGA in terms of the platform. Although the GPU has higher peak performance and memory bandwidth, our FPGA implementation achieves a latency reduction of 2.78×2.78$\times$2.78 × and a throughput improvement of 2.1×2.1$\times$2.1 ×. This speedup is attributed to the optimizations mentioned above adopted in our implementation.
@@
6 Conclusion
This work presented a novel architecture based on MLPs and graph convolutional layers. We benchmarked our model on popular grayscale image datasets and highlighted the strength of our model in terms of its experimental performance and model complexity metrics. We emphasized the importance of a lightweight image classifier and examined the vast domain of grayscale image datasets. Overall, a lightweight and low-latency image classifier can improve the efficiency and effectiveness of various applications that rely on image processing.
Our model performs well in various domains, including general image classification, SAR ATR, and medical image classification. Additionally, we implement our model on FPGA to show that our model is hardware-friendly, which is an added benefit of any image classification model.
We presented the novelties of our model, which include viewing an image as a vector in a fully connected manner and inputting the resultant into a single-layer GCN. We also found that graph neural networks can increase the accuracy and reduce the variance in performance of shallow models like ours. Additionally, a simple batch-wise attention term allows our shallow model to capture interdependencies between images and form connections for classification. These novelties result in a highly low latency and high throughput model.

### v4

@@ 1 Introduction
  • We implement our model on a state-of-the-art FPGA board, Xilinx Alveo U200. Compared with state-of-the-art GPU implementation, our FPGA implementation achieves a 2.78×2.78$\times$2.78 × speedup in latency and a 2.1×2.1$\times$2.1 × speedup in throughput.
@@ 3.1 MLP Approaches
Our model combines various components of simple models and is inherently different from current works in low-latency image classification. Some recent architectures involve simple MLP-based models. Touvron et al. introduced ResMLP [2], an image classifier based solely on MLPs. ResMLP is trained in a self-supervised manner with patches interacting together. Touvron et al. highlight their model’s high throughput properties and accuracy. ResMLP uses patches from the image and alternates linear layers where patches interact and a two-layer feed-forward network where channels interact independently per patch. Additionally, MLP-Mixer [3] uses a similar patching method, which also attains competitive accuracy on RGB image datasets compared to other CNNs and transformer models. Our proposed method uses the results from a single-layer MLP to feed into a graph neural network, during which we skip the information from the three-channel RGB setting and only consider the single-channel grayscale problem. This is inherently different from the methods mentioned earlier, as they use patching approaches while we focus on the vectorization of pixels.
@@ 3.2 Graph Image Construction Methods
Han et al. [7] form a graph from the image by splitting the image into patches, much like a transformer. A deep graph neural network learns on the patches similarly to a transformer but in a graph structure. Our structure does not form a graph where each patch is a node in a graph; instead, we create a graph from the resultant of a vectorized image passed through a fully connected layer.
@@ 4 Overview and Architecture
Our model vectorizes a batch of images, allowing us to use a fully connected layer pixel-wise for low computation time rather than relying on convolutional neural networks. The proposed method uses the resultant from FCs and constructs a graph batch-wise, where each node corresponds to the flattened image outputted from an FC. We then apply a batch-wise attention term, which is inputted into an FC for classification. [^1] We make our code publicly available at https://github.com/GECCOProject/GECCO
Image Vectorization. For each image in a batch, we view the image as a vector. For a tensor $\mathbf{X}\in\mathbb{R}^{B\times H\times W}$ where $B$ is the batch size, $H$ and $W$ are the height and width of an image; we flatten the tensor to $\mathbf{X}_{1}\in\mathbb{R}^{B\times(H\cdot W)}$. Viewing an image as a vector allows our model to skip the traditional convolutional neural network, which views the image as a grid and cuts computation time.
@@
Fully Connected Layer. We input $\mathbf{X}_{1}$ into a fully connected layer (FC) with output dimensionality $D_{out}$. Formally, $\mathbf{X}_{2}=\sigma(\mathbf{X}_{1}\mathbf{W}_{1}+\mathbf{b}_{1})$, where $\sigma$ is the ReLU function, $\mathbf{W}_{1}$ is a learned weight matrix, $\mathbf{b}_{1}$ is a bias term, and $\mathbf{X}_{2}\in\mathbb{R}^{B\times D_{out}}$.
After the fully connected layer, we apply a dropout layer and the ReLU function to $\mathbf{X}_{2}$, yielding $\mathbf{X}_{3}$, such that the resultant dimensionality of $\mathbf{X}_{3}$ is $\mathbb{R}^{B\times D_{out}}$.
@@
Graph Construction. We construct a graph batch-wise from $\mathbf{X}_{3}$. This means that for each batch, a vectorized image is each node in the graph with feature size $\mathbb{R}^{D_{out}}$, and each image is connected to every other image in a batch. Formally, we calculate the adjacency matrix $\mathbf{A}$ as $\mathbf{A}_{ij}=1$, which connects all nodes.
@@
Graph Convolution. Our single graph convolutional layer learns from similar features of images within its minibatch. Generally, a graph convolutional layer updates the representations of nodes by aggregating each node’s neighbor’s representation. We can write a graph convolutional layer as $\bm{h^{\prime}}_{i}=f_{\theta}\left(\bm{h}_{i},\text{\scriptsize{AGGREGATE}}\left(\{\bm{h_{j}}\mid j\in\mathcal{N}_{i}\}\right)\right)$. In our case, the input for each node $\bm{h_{i}}$ is the output from each vector in $\mathbf{X}_{3}$.
Applying graph convolution to $\mathbf{X}_{3}$ yields $\mathbf{X}_{4}$. Formally, $\mathbf{X}_{4}=\sigma\left(\mathbf{A}\mathbf{X}_{3}\mathbf{W}_{2}\right)$, where $\mathbf{W}_{2}$ is a learned weight matrix. After the graph convolution, we apply batch normalization and max-pooling operations to $\mathbf{X}_{4}$, resulting in a dimensionality of $\mathbb{R}^{B\times\lfloor D_{out}/2\rfloor}$.
@@
Batch-wise Attention, Residual Connections, & Output. We propose a batch-wise attention term defined as
@@
where $\sigma$ is the sigmoid function. This term allows the model to capture similar features from each image to another batch-wise.
The residual connection is defined $\mathbf{X}_{6}=\mathbf{X}_{5}+\mathbf{X}_{4}$. The residual term makes the learning process easier and more stable. By multiplying a softmax-like term with the output of a previous graph convolution, we weigh the correspondence of each pixel compared to other similar pixels within similar images batch-wise. We then feed the residual term into an FC inputted into the softmax function for classification results.
Model Structure Discussion
We justify our model’s design choices by considering the following theoretical aspects.
  1. 1.
    The batch-wise attention term allows the model to further capture similar features from each image to another batch-wise. Relating similar properties from images to each other boosts accuracy in our case of a shallow model. Additionally, our batch-wise attention term is similar in spirit to the mechanism proposed by [9], which allows the model to capture long-range dependencies across the entire image.
  2. 2.
    The batch-size hyperparameter is crucial in our model. A larger batch size allows the model to capture more dependencies across images, which is crucial for understanding complex image patterns. We refer to the work of [10] for a detailed analysis of the impact of batch size on the performance of GNNs.
  3. 3.
    If the batch size for a given dataset is $1$, the model eliminates the graph construction phase, making the term $\mathbf{X}_{3}$ fed directly into the FC and softmax for classification.
  4. 4.
    The residual connection term makes the learning process easier and more stable. we refer to [11] for a more detailed analysis of the impact of residual connections on shallow models.
@@ 5.1 Datasets
  • MSTAR is a SAR ATR dataset with a training size of $2747$ and testing size of $2425$ SAR images of $10$ different vehicle categories. We resize each image in the dataset to $(128,128)$ pixels.
@@ 5.2.1 Backbone
GCN [15] $97.12\%$ $54.58\pm 8.47$ $4.68\pm 0.72$
TAGConv [16] $97.36\%$ $58.39\pm 9.01$ $4.37\pm 0.67$
SAGEConv [17] $97.72\%$ $58.84\pm 8.98$ $4.32\pm 0.66$
ChebConv [18] $97.79\%$ $51.98\pm 8.65$ $4.91\pm 0.82$
ResConv [19] $\bf{97.84\%}$ $\bf{62.44\pm 9.87}$ $\bf{4.10\pm 0.65}$
We use the following hyperparameters for our experiments.
Table 2: Feature Lengths, Optimizer, and Batch Size for Each Dataset
Dataset Feature Length Optimizer Batch Size
MNIST $64$ Adam $64$
Fashion-MNIST $64$ Adam $64$
MSTAR $86$ Adam $64$
CXR $112$ Adam $64$
@@ 5.2.2 Experimental Performance
In Tables 3, 4, 5, and 6, we present a summary of our findings. We report the best-performing accuracy, average throughputs, and latencies with their standard deviations. Our model outperforms every other model in terms of throughput and latency across all datasets, leads accuracy on the MSTAR dataset, and performs competitively in terms of accuracy on all datasets.
@@ 5.2.3 Model Complexity Metrics
Model complexity metrics for this paper include the number of multiply-accumulate operations (MACs), the number of model parameters, the model size, and the number of layers. In other words, suppose accumulator $a$ counts an operation of arbitrary $b,c\in\mathbb{R}$. We count the number of multiply-accumulate operations as $a\leftarrow a+(b\times c)$. Additionally, the layer count metric is an essential factor of latency. Increasing the number of layers will also improve the latency of a model’s inference. The goal of an effective machine learning model is to maximize throughput while minimizing the number of MACs and the number of layers, in our case.
@@
GECCO $\bf{5.89\times 10^{6}}$ $\bf{1.90\times 10^{4}}$ $\bf{0.075}$ $\bf{10}$
@@ 5.3 Discussion
Across multiple datasets, GECCO achieves leading or competitive accuracy compared to other state-of-the-art image classifiers. GECCO outperforms other machine learning models regarding model complexity, highlighting our model’s low latency and lightweight properties.
It is difficult for our model to generalize to the RGB setting. We attribute this challenge to the vectorization process that our model uses. Learning on three channels poses a complexity challenge, as GECCO is very shallow and simple, thus making it challenging to learn on three separate channels. Additionally, our model is optimized for a low-complexity dataset regime, as they are much too complex for our shallow model. Datasets like grayscale ImageNet or CIFAR variants are unsuitable for GECCO as they pose a complexity challenge.
Our proposed method does not make use of positional embeddings or class tokens. GECCO can learn essential features using the weighted residual term. Additionally, we tested the addition of positional embeddings and class tokens and found no improvement in accuracy across various datasets. We note that the $\mathbf{X}_{5}$ residual term adds positional awareness to the model.
@@ 5.4 FPGA Implementation
We develop an accelerator for the proposed model on a state-of-the-art FPGA, Xilinx Alveo U200 board, to further highlight the model’s efficiency and compatibility with hardware. It has 3 Super Logic Regions (SLRs), 4 DDR memory banks, 1182k Look-up tables, 6840 DSPs, 75.9 Mb of BRAM, and 270 Mb of URAM. The FPGA kernels are developed using the Xilinx High-level Synthesis (HLS) tool to expedite the design process.
We perform the following optimizations for our FPGA design: (1) Portable design: We design a parameterized hardware template using HLS. It is portable to different FPGA platforms, including embedded and data-center FPGAs. (2) Resource sharing: The model is executed layer-by-layer. Each layer in the model is decomposed into basic kernel functions. The basic kernel functions, including matrix multiplication, elementwise activation, column-wise and row-wise summations, max pooling, and various other elementwise operations, are implemented separately and subsequently invoked within their corresponding layers. Due to the reuse of these fundamental kernel functions across multiple layers, FPGA resources are shared among the different layers, maximizing runtime hardware resource utilization. (3) One-load strategy: We employ a one-time data load strategy to load the required data from DDR only once. All other data required for the computations are stored in on-chip memory, reducing inference latency. Figure 3 illustrates the overall hardware architecture of our design.
@@
We compare our FPGA implementation with the baseline GPU implementation. The GPU baseline is executed on an NVIDIA RTX A5000 GPU, which operates at 1170 MHz and has a memory bandwidth of 768 GB/s. On the other hand, the FPGA operates at 300 MHz and has an external memory bandwidth of 77 GB/s. The GPU baseline is comparable with the FPGA in terms of the platform. Although the GPU has higher peak performance and memory bandwidth, our FPGA implementation achieves a latency reduction of 2.78×2.78$\times$2.78 × and a throughput improvement of 2.1×2.1$\times$2.1 ×. This speedup is attributed to the optimizations mentioned above that were adopted in our implementation.
@@
6 Conclusion & Future Work
This work introduced a novel architecture combining MLPs and graph convolutional layers, benchmarked on popular grayscale image datasets. The model demonstrated strong performance and low complexity, highlighting the importance of lightweight, low-latency image classifiers for various applications. Its efficacy was shown across general image classification, SAR ATR, and medical image classification, with an FPGA implementation underscoring its hardware compatibility. Key innovations include using a single-layer GCN, which, along with batch-wise attention, enhances accuracy and reduces variance. Future work should explore extending this approach to color image datasets and other domains, optimizing the architecture for even greater efficiency, and further investigating the potential of graph neural networks in shallow models.
7 Acknowledgement
This work is supported by the DEVCOM Army Research Lab (ARL) under grant W911NF2220159. Distribution Statement A: Approved for public release. Distribution is unlimited.